OpenStudy (anonymous):

Find the greatest number that will divide 43, 91 and 183 so as to leave the same remainder in each case. A. 4 B. 7 C. 9 D. 13

OpenStudy (anonymous):

4

OpenStudy (anonymous):

plz explain

OpenStudy (anonymous):

just plug in all the values and check..

OpenStudy (anonymous):

hello him1618 ...plz explain the step

myininaya (myininaya):

one sec let me see if i can come up with a good explanation besides just plugging them in which works too please hold

myininaya (myininaya):

the way him was talking about might be the only way

OpenStudy (anonymous):

k

OpenStudy (anonymous):

how to plug in all the values?

myininaya (myininaya):

we are looking for the same remainder we want to chose from the possible list of divisors A,B,C,D that will give us this when we divide 43,91,183 183/4=45+138+3/4 (3 is the remainder) 91/4=22+3/4 ( 3 is the remainder) 43/4=5+3/4 ( 3 is the remainder) so we are done we do not have to go through B,C, and D since we have our answer

OpenStudy (anonymous):

ohhh tq .. :)

myininaya (myininaya):

np